OpenStudy (anonymous):

how would i solve this. b+8 - b+4? then b+4 - b ?

OpenStudy (mathstudent55):

Combine like terms. Like terms are terms that have the same variable part. All the terms with b are like terms. All terms without b are like terms. b + 8 - b + 4 = b - b + 8 + 4 = 12

OpenStudy (anonymous):

I see. so for b+4 -b i'd do , b-b+4 =4?

OpenStudy (mathstudent55):

correct

OpenStudy (mathstudent55):

For x + y + 6 - x + y +9, you'd do x - x + y + y + 6 + 9 = 2y + 15

OpenStudy (anonymous):

thank you.

OpenStudy (mathstudent55):

wlcm